Zeronknight (zeronknight):

What number is 30% of 40? * 13.3 30 12 120

Vocaloid (vocaloid):

percent means per 100 so 30 percent means 30 per 100, or (30/100) so we just multiply (30/100) * 40

Zeronknight (zeronknight):

That doesn't exactly make much since. 30/100 x 40?

Vocaloid (vocaloid):

yes what part doesn't make sense?

sammixboo (sammixboo):

We need to set up a proportion

sammixboo (sammixboo):

To better understand this problem

sammixboo (sammixboo):

\(\tt \dfrac{x}{40} = \dfrac{30}{100}\)

Zeronknight (zeronknight):

30 x 40?

OpenStudy (skullpatrol):

This is the basic formula: $$\Huge x\%=\dfrac{a}{b}$$

OpenStudy (skullpatrol):

have you seen that^ before?

sammixboo (sammixboo):

Yes, then divide that by 100

OpenStudy (abdullahm):

As I like to say: whenever you see 'of' think of that as multiplication whenever you see 'is' think of that an equal sign and change 'what number' to a variable like x What number is 30% of 40? * x = 30% * 40 % like I told you in your previous post is over 100 so x = (30/100) * 40

OpenStudy (skullpatrol):